OpenStudy (anonymous):

If D varies directly with t and D is 105 when t is 3, find D when t is 8. 35 280 840

OpenStudy (alexandervonhumboldt2):

3 =>105 8 =>x x/8=105/3 x=35*8=280

OpenStudy (alexandervonhumboldt2):

use ratios